# Retrieval-augmented Few-shot Medical Image Segmentation with Foundation Models

### v1

@@ Abstract
Medical image segmentation is crucial for clinical decision-making, but the scarcity of annotated data presents significant challenges. Few-shot segmentation (FSS) methods show promise but often require retraining on the target domain and struggle to generalize across different modalities. Similarly, adapting foundation models like the Segment Anything Model (SAM) for medical imaging has limitations, including the need for finetuning and domain-specific adaptation. To address these issues, we propose a novel method that adapts DINOv2 and Segment Anything Model 2 (SAM 2) for retrieval-augmented few-shot medical image segmentation. Our approach uses DINOv2’s feature as query to retrieve similar samples from limited annotated data, which are then encoded as memories and stored in memory bank. With the memory attention mechanism of SAM 2, the model leverages these memories as conditions to generate accurate segmentation of the target image. We evaluated our framework on three medical image segmentation tasks, demonstrating superior performance and generalizability across various modalities without the need for any retraining or finetuning. Overall, this method offers a practical and effective solution for few-shot medical image segmentation and holds significant potential as a valuable annotation tool in clinical applications.
@@ I Introduction
Medical image segmentation plays a critical role in diagnosis, treatment planning, and the quantification of tissue volumes [1, 2]. Accurate segmentation enables clinicians to precisely delineate anatomical structures and abnormalities within medical images, which is essential for making informed clinical decisions. However, unlike natural image segmentation, medical image segmentation faces unique challenges due to the limited availability of data and annotations [3, 4, 5]. Obtaining precise annotations requires the expertise of trained medical professionals and is time-consuming, making it difficult to gather enough labeled samples for supervised training. In practice, while acquiring a large number of annotations is often impractical, obtaining a smaller set, such as a few dozen, is more feasible. Maximizing the utility of these limited samples to create reliable segmentation masks without compromising the accuracy is a pressing challenge in the field.
@@
Few-shot segmentation (FSS) offers promising potential to address the aforementioned challenges in the medical imaging domain, particularly the scarcity of labeled data [6]. Various approaches have been developed to enable models to learn from a limited number of annotated samples, such as ALPNet [7], ADNet [8],Q-Net [9], and CAT-Net [10]. Despite their effectiveness and success, these methods have certain limitations. They often require retraining on a target domain with known classes before being applied to new classes within the same domain, limiting their ability to generalize across different medical image modalities. Additionally, most of these methods employ a 1-way 1-shot setting, where neighboring slices are used as support images for FSS. While the 1-shot setting can be valuable in research, acquiring a small number of annotated samples is feasible and aligns better with clinical practices. Furthermore, reliance on neighboring slices can constrain the model’s flexibility and effectiveness, especially when annotations for neighboring slices are unavailable for new subjects or when there is significant anatomical variability between slices. These limitations restrict the practical application and effectiveness of these models in diverse clinical scenarios.
@@
Building on this intuition, we propose a novel framework that adapts DINOv2 [20] and SAM 2 [17] for few-shot medical image segmentation. The framework features a retrieval module that uses DINOv2 to generate embeddings and retrieve similar samples from limited annotated data. The segmentation network, adapted from SAM 2, encodes these retrieved images and masks as memories stored in a memory bank. The memory attention mechanism then integrates relevant features and anatomical structures from these stored memories, enhancing the input image features for the decoder which generates the final segmentation mask with greater precision and accuracy. Notably, our framework directly utilizes the structure and pre-trained weights without any retraining or finetuning in the target domain. When applied to a new domain, only the samples in the retrieval database need to be updated. We evaluate the proposed framework on three medical image segmentation datasets, including ACDC, CMR T1-Map, and Fluoroscopy Image datasets. Building on these foundation models, our method demonstrates superior performance and generalizes well across different modalities and contrasts without any retraining or finetuning. Our contributions in this work are summarized as follows:
  • We introduce a novel framework that integrates retrieval augmentation and memory-driven approach, leveraging foundation models like DINOv2 and SAM 2 for accurate medical image segmentation with limited annotated data.
@@ II-C Few-shot Medical Image Segmentation
Few-shot segmentation (FSS) has been widely explored in the medical imaging field, where data and annotations are usually scarce. Typically, FSS frameworks follow the Prototype Alignment Network (PANet) approach, which learns class-specific prototype representations from a few support images and then performs segmentation on the query images by matching each pixel to the learned prototypes [6]. For the FSS in medical imaging domain, ALPNet enhances PANet by introducing superpixel-based self-supervised learning, eliminating the need for manual annotations, and by incorporating an adaptive local prototype pooling module to address the foreground-background imbalance in medical image segmentation [7]. The Anomaly Detection-inspired Network (ADNet) takes a different approach by relying on a single foreground prototype to compute anomaly scores for all query pixels, with segmentation achieved by thresholding these scores using a learned threshold [8]. Q-Net improves the ADNet by adding a query-informed threshold adaptation module and a query-informed prototype refinement module [9]. CAT-Net, which leverages a cross-masked attention transformer, enhances the FSS by mining the correlations between the support and query images and restricts the focus to relevant foreground information [10]. Despite the success of these FSS methods, they still require further retraining on the target domain and are typically trained and applied within the same domain for few-shot segmentation of new classes. This limitation significantly restricts their practicality in real-world applications. Our approach leverage the foundation models to perform the FSS without retraining in new domain, enabling more flexible and efficient segmentation across diverse medical imaging applications.
@@ IV-A2 CMR T1-Map Dataset
The CMR T1-Map dataset is dedicated to segmenting the myocardium (Myo) in T1 mapping CMR images, acquired under varying scanning protocols. To create a representative dataset, we randomly selected 3 subjects from a pool of 163 patients, extracting a total of 50 axial slices for training. Similarly, an additional 3 subjects with 50 slices were selected for the test dataset.
@@ IV-A3 Fluoroscopy Image Dataset
The Fluoroscopy Image dataset is designed for the segmentation of the left coronary artery (LCA). It includes data from 9 subjects with annotations, with 5 subjects used to construct the retrieval database and 4 subjects designated for the test dataset. For each subject, we selected 10 consecutive frames from two different positional angles, resulting in a total of 20 frames per subject. This setup yields 100 frames for the retrieval database and 80 frames for the test dataset.
@@ IV-C Segmentation Results on Three Datasets
SAM 2 (1 Pos Point) 0.4146 0.4565 0.6612 0.4483 0.3761
SAM 2 (2 Pos Points) 0.4268 0.4604 0.6610 0.4436 0.3849
SAM 2 (1 Pos&Neg Points) 0.4441 0.4794 0.6952 0.4215 0.3844
SAM 2 (2 Pos&Neg Points) 0.4133 0.5004 0.6801 0.4357 0.4858
Ours 0.6729 0.7757 0.8472 0.8238 0.8431
@@ IV-D Comparison with Supervised Methods
In this subsection, we compare the segmentation performance of our method with fully-supervised methods on the ACDC dataset, as presented in Table Table II. The table highlights the results of traditional fully-supervised models like U-Net [15] and SwinUNETR [37] under both full and limited data settings. When trained with the full dataset of over 1800 samples, U-Net and SwinUNETR achieve high dice coefficients. However, in the limited sample setting with only 50 samples, the performance of these fully-supervised models drops significantly, with U-Net and SwinUNETR showing markedly lower DSCs, especially for the right ventricle (RV) and myocardium (Myo).
TABLE II: Comparison of segmentation performance on the ACDC dataset between fully-supervised methods and our approach.
Setting Methods ACDC
RV Myo LV
Full Data (1808 slices) U-Net 0.8743 0.8785 0.9473
SwinUNETR 0.7623 0.8346 0.9198
Limited Data (50 slices) U-Net 0.1996 0.4794 0.5938
SwinUNETR 0.2792 0.4202 0.4542
Ours 0.6729 0.7757 0.8472
In contrast, our method, which leverages foundation models to effectively utilize limited data, achieves much higher dice coefficients than supervised model. Remarkably, even when compared to the supervised methods trained on the full dataset, our method’s performance is only about 0.1 DSC behind, suggesting its robustness and adaptability in scenarios with limited annotated data. This close performance gap demonstrates the effectiveness of our approach, particularly in medical image segmentation where acquiring large annotated datasets can be challenging.
IV-E Ablation Studies
@@
Table III compares the segmentation performance on the ACDC dataset using two retrieval strategies: Random and DINOv2, with two different numbers of queried images, #8 and #16.
TABLE III: Comparison of segmentation performance on the ACDC dataset using different retrieval strategies (Random vs. DINOv2) and varying numbers of queried images (#8 vs. #16).
@@ V-A Annotation Tools
V-B Limitations
@@
Figure 5 illustrates the segmentation performance of our method on challenging basal and apical slices from the ACDC dataset. Despite our method’s overall effectiveness demonstrated in previous sections, the segmentation results for these particular slices are suboptimal. This is a common challenge even for supervised methods, as the basal and apical slices often present more anatomical variations and smaller regions of interest, making accurate segmentation difficult. The queried images visualized indicate that the retrieved samples, while semantically similar, may not provide adequate support due to the small size of the target regions and the inherent heterogeneity in medical images. This suggests that the retrieval process, which uses whole image features, might not always capture the fine details needed for precise segmentation of small objects. Furthermore, the limited size and variability of the database mean that sufficiently similar images to support accurate segmentation might not be available. These factors underscore the ongoing challenges in medical image segmentation, particularly for small and variable anatomical structures.

### v2

@@ Abstract
Medical image segmentation is crucial for clinical decision-making, but the scarcity of annotated data presents significant challenges. Few-shot segmentation (FSS) methods show promise but often require training on the target domain and struggle to generalize across different modalities. Similarly, adapting foundation models like the Segment Anything Model (SAM) for medical imaging has limitations, including the need for finetuning and domain-specific adaptation. To address these issues, we propose a novel method that adapts DINOv2 and Segment Anything Model 2 (SAM 2) for retrieval-augmented few-shot medical image segmentation. Our approach uses DINOv2’s feature as query to retrieve similar samples from limited annotated data, which are then encoded as memories and stored in memory bank. With the memory attention mechanism of SAM 2, the model leverages these memories as conditions to generate accurate segmentation of the target image. We evaluated our framework on three medical image segmentation tasks, demonstrating superior performance and generalizability across various modalities without the need for any retraining or finetuning. Overall, this method offers a practical and effective solution for few-shot medical image segmentation and holds significant potential as a valuable annotation tool in clinical applications.
@@ I Introduction
Medical image segmentation plays a critical role in diagnosis, treatment planning,enabling precise analysis for various clinical applications. [1, 2, 3]. Accurate segmentation enables clinicians to precisely delineate anatomical structures and abnormalities within images, which is essential for making informed clinical decisions. However, unlike natural image segmentation, medical image segmentation faces unique challenges due to the limited availability of data and annotations [4, 5, 6]. Obtaining precise annotations requires the expertise of trained medical professionals and is time-consuming and labor-intensive, making it difficult to gather enough labeled samples for supervised training. In practice, while acquiring a large number of annotations is often impractical, obtaining a smaller set, such as a few dozen, is more feasible. Maximizing the utility of these limited samples to create reliable segmentation masks without compromising the accuracy is a pressing challenge in the field.
@@
Few-shot segmentation (FSS) offers promising potential to address the aforementioned challenges in the medical imaging domain, particularly the scarcity of labeled data [7]. Various approaches have been developed to enable models to learn from a limited number of annotated samples, such as ALPNet [8], ADNet [9],Q-Net [10], and CAT-Net [11]. Despite their effectiveness and success, these methods have certain limitations. They often require self-supervised pretraining on the target domain [8, 9], or training with known classes before being applied to new classes [11], limiting their ability to generalize across different image modalities. Additionally, most of these methods employ a 1-way 1-shot setting. While this setting can be valuable in research, acquiring a small number of annotated samples is feasible and aligns better with clinical practices. Furthermore, the compromised performance of these methods may not match that of supervised methods trained on large datasets, limiting their ability to reduce the annotation burden. These limitations restrict the practical application and effectiveness of these models in diverse clinical scenarios.
@@
Building on this intuition, we propose a novel framework that adapts DINOv2 [21] and SAM 2 [18] for few-shot medical image segmentation. The framework introduces a retrieval module that uses DINOv2 to generate embeddings and retrieve similar samples from limited annotated data. The segmentation network, adapted from SAM 2, encodes these retrieved images and masks as memories stored in a memory bank. The memory attention mechanism then integrates relevant features and anatomical structures from these stored memories, enhancing the input image features for the decoder which generates the final segmentation mask with greater precision and accuracy. Notably, our framework directly utilizes the structure and pre-trained weights without any retraining or finetuning in the target domain. When applied to a new domain, only the samples in the retrieval database need to be updated. We evaluate the proposed framework on three medical image segmentation datasets, including public ACDC dataset [22], and in-house CMR T1-Map, and Fluoroscopy Image datasets. Building on these foundation models, our method demonstrates superior performance and generalizes well across different modalities and contrasts. Compared with other few-shot medical image segmentation methods, our method achieve state-of-the-art performance without any retraining or fine-tuning in ABD-MRI dataset [23]. Our contributions in this work are summarized as follows:
  • We introduce a novel framework that integrates retrieval augmentation and memory-driven approach, leveraging foundation models like DINOv2 and SAM 2 for accurate few-shot medical image segmentation with limited annotated data.
@@ II-C Few-shot Medical Image Segmentation
Few-shot segmentation (FSS) has been widely explored in the medical imaging field, where data and annotations are usually scarce. Typically, FSS frameworks follow the Prototype Alignment Network (PANet) approach, which learns class-specific prototype representations from a few support images and then performs segmentation on the query images by matching each pixel to the learned prototypes [7]. For the FSS in medical imaging domain, ALPNet enhances PANet by introducing superpixel-based self-supervised learning, eliminating the need for manual annotations, and by incorporating an adaptive local prototype pooling module to address the foreground-background imbalance in medical image segmentation [8]. The Anomaly Detection-inspired Network (ADNet) takes a different approach by relying on a single foreground prototype to compute anomaly scores for all query pixels, with segmentation achieved by thresholding these scores using a learned threshold [9]. Q-Net improves the ADNet by adding a query-informed threshold adaptation module and a query-informed prototype refinement module [10]. CAT-Net, which leverages a cross-masked attention transformer, enhances the FSS by mining the correlations between the support and query images and restricts the focus to relevant foreground information [11]. Despite the success of these FSS methods, they still require self-supervised pretraining or training on the base classes before applying to new classes in the same domain. This limitation restricts their practicality in real-world applications. Our approach leverage the foundation models to perform the FSS without any training in new domain, enabling more flexible and efficient segmentation while achieving state-of-the-art performance.
@@ IV-A2 CMR T1-Map Dataset
The CMR T1-Map dataset is dedicated to segmenting the myocardium (Myo) in T1 mapping CMR images, acquired under varying scanning protocols. To create a representative dataset, we randomly selected 3 subjects from a pool of 163 patients, extracting a total of 50 axial slices for retrieval. Similarly, an additional 3 subjects with 50 slices were selected for the test dataset. For comparison, the supervised training uses the same test dataset and a training dataset with an 402 axial slices from 25 subjects.
@@ IV-A3 Fluoroscopy Image Dataset
The Fluoroscopy Image dataset is designed for the segmentation of the left coronary artery (LCA). It includes data from 9 subjects with annotations, with 5 subjects used to construct the retrieval database and 4 subjects designated for the test dataset. For each subject, we selected 10 consecutive frames from two different positional angles, resulting in a total of 20 frames per subject. This setup yields 100 frames for the retrieval database and 80 frames for the test dataset. For supervised methods, we sampled an additional 80 frames for each subject, resulting in a training set of 500 frames.
@@ IV-C Segmentation Results on Three Datasets
SAM 2 (1 Pos Point) 0.4146 0.4565 0.6612 0.4483 0.2788
SAM 2 (2 Pos Points) 0.4268 0.4604 0.6610 0.4436 0.2736
SAM 2 (1 Pos&Neg Points) 0.4441 0.4794 0.6952 0.4215 0.2592
SAM 2 (2 Pos&Neg Points) 0.4133 0.5004 0.6801 0.4357 0.3477
Ours 0.6729 0.7757 0.8472 0.8238 0.7029
@@ IV-D Comparison with Supervised Methods
We compare the segmentation performance of our method with fully-supervised methods such as U-Net [16] and SwinUNETR [40] on three datasets, and present the results under both full and limited data settings in Table II. When trained with full dataset, U-Net and SwinUNETR achieve high DSCs in ACDC and CMR T1-map datasets. However, in the limited sample setting with only 50 samples, the performance of these fully-supervised models drops, with U-Net and SwinUNETR showing markedly lower DSCs, especially for the right ventricle (RV) and myocardium (Myo) in ADCD dataset.
TABLE II: Comparison of segmentation performance on the ACDC, CMR T1-MAP, and Fluoroscopy Image datasets between fully-supervised methods and our approach.
Settings Methods ACDC CMR T1-Map Fluoroscopy Image
RV Myo LV Myo
Full Data U-Net 0.8743 0.8785 0.9473 0.8700 0.6539
SwinUNETR 0.7623 0.8346 0.9198 0.8769 0.6058
Limited Data (50 slices) U-Net 0.1996 0.4794 0.5938 0.7792 0.5760
SwinUNETR 0.2792 0.4202 0.4542 0.8147 0.5558
Ours 0.6729 0.7757 0.8472 0.8238 0.7029
In contrast, our method, which leverages the foundation model, achieves higher DSCs than supervised model in the limited data setting. Notably, on the Fluoroscopy image datasets, our method with limited data outperforms the supervised methods trained on the full dataset by 5%-10%. For the ACDC and CMR T1-Map datasets, our method’s performance is only about 10% lower in DSC. This suggests that the segmentation priors learned by the foundation model SAM 2 are highly generalizable. With a small amount of data, the model can quickly adapt to different tasks and achieve comparable or even superior performance compared to supervised methods designed for specific segmentation tasks.
IV-E Comparison with Other Few-shot Medical Image Segmentation Methods
We compare our method with other state-of-the-art (SOTA) methods for few-shot medical image segmentation, including PA-Net [7], ALP-Net [8], AD-Net [9], Q-Net [10] and CAT-Net [11]. For this comparison, we use the Abd-MRI dataset [23], which consists of 20 abdominal MRI scans with annotations for the left kidney (LK), right kidney (RK), spleen (Spl), and liver (Liv). We follow the 1-way 1-shot scenario and experiment setting described in [11] with 5-fold cross-validation. However, since our method does not require any training, we directly apply it to the test set for each fold. The results are demonstrated in Table III.
TABLE III: Comparison of segmentation performance on the Abd-MRI dataset [23] between our method and other few-shot medical image segmentation methods.
Methods LK RK Spl. Liv. Avg.
PA-Net [7] 0.4771 0.4795 0.5873 0.6499 0.5485
ALP-Net [8] 0.7363 0.7839 0.6702 0.7305 0.7302
AD-Net [9] 0.7189 0.7602 0.6584 0.7603 0.7270
Q-Net [10] 0.7405 0.7752 0.6743 0.7871 0.7443
CAT-Net [11] 0.7401 0.7890 0.6883 0.7898 0.7518
Ours 0.7779 0.8581 0.7586 0.8793 0.8185
Our method achieves the highest DSCs across all evaluated organs. Specifically, our method obtains DSC of 0.7779 for the left kidney (LK), 0.8581 for the right kidney (RK), 0.7586 for the spleen (Spl.), and 0.8793 for the liver (Liv.), resulting in an average DSC of 0.8185. This outperforms the best-performing SOTA method, CAT-Net, which achieves an average DSC of 0.7518. The improvement is particularly notable for the spleen and liver, where our method surpasses CAT-Net by 7.03% and 8.95%, respectively. These results demonstrate the superior generalizability and adaptability of our method in few-shot medical image segmentation tasks. Notably, unlike those dedicated few-shot segmentation approaches, our method does not require any pretraining or training on base classes before testing on novel classes within the same data modality. This flexibility makes it well-suited for real-world clinical applications, where annotated data is often scarce, but the target regions and data modalities are diverse.
IV-F Ablation Studies
@@
Table IV compares the segmentation performance on the ACDC dataset using two retrieval strategies: Random and DINOv2, with two different numbers of queried images, #8 and #16.
TABLE IV: Comparison of segmentation performance on the ACDC dataset using different retrieval strategies (Random vs. DINOv2) and varying numbers of queried images (#8 vs. #16).
@@ V-A Annotation Tools
V-B Extend to 3D Medical Image Segmentation
Our current method can be directly extend to 3D medical image segmentation by treating the 3D volume as a series of 2D slices. The database can store all slices from 3D volume of a subject or multiple subjects, and when a new subject is introduced, the model can segment the 3D volume slice by slice, retrieving the most similar slices from the database to guide the segmentation. For future work, performance could be further enhanced by adapting SAM 2 for 3D segmentation and incorporating specialized adaptors to capture volumetric information and spatial dependencies across slices.
V-C Running Time Analysis
We have conducted a runtime comparison to assess the computational cost of the retrieval and memory-attention steps. Specifically, the DINOv2 encoding and retrieval process in our method takes an averaged 8.78 ms per slice with one NVIDIA A100 GPU, and the complete segmentation using our method requires 52.96/170.94/515.2 ms per slice when having 1/4/16 retrieved images to guide the segmentation. In comparison, SAM 2 with 4 point prompts takes 9.987 ms per slice. Although the increased runtime may limit real-time performance for high-throughput batch processing, it remains practical for clinical scenarios where accuracy and adaptability to limited data are more critical than processing speed. While SAM 2 with point prompts requires less time, its segmentation performance is not sufficient for clinical applications or to effectively reduce the annotation burden. In contrast, the segmentation results from our few-shot method can be used to train supervised segmentation models, helping to reduce the need for manual annotations or potentially being adapted for real-time processing with further optimization.
V-D Influence of Database Size
We expand the database to include the full ACDC dataset (1808 images) for retrieval to invesitage the influence of including more samples in the database. We found that as the database size increases, the performance gain is minimal. The DSC for the right ventricle (RV) changed slightly from 0.6729 to 0.6695, for the myocardium (Myo) from 0.7757 to 0.7772, and for the left ventricle (LV) from 0.8472 to 0.8693. These results suggest that our model does not directly benefit from a larger database size, indicating that simply adding more diverse samples may not be the key to improving performance.
To further investigate the factors driving model performance, we conducted an experiment where we included the ground truth mask and the image to be segmented into the memory bank. This led to a remarkable improvement: the DSC for the right ventricle (RV) increased from 0.6729 to 0.9281, for the myocardium (Myo) from 0.7757 to 0.9075, and for the left ventricle (LV) from 0.8472 to 0.9577. This indicates that the similarity between the queried images and the test image is crucial for improving segmentation performance, as more similar images can provide more relevant and contextually aligned information for the model.
V-E Limitations
FOV Variations. In our experiments, the test image and the retrieved query images from the database have similar FOV and content, which enables DINOv2 to effectively measure global similarity. However, a significant FOV mismatch between the test and query images could limit the proposed method’s applicability in annotation tools, such as when the database contains a full-body MRI image while the test image only contains a specific region like the heart. A potential solution could be to leverage the ground truth segmentation masks of full-body MRI images to crop them into smaller patches of varying sizes, each containing a specific organ or a combination of organs. This would increase the diversity of the database and enhance the model’s ability to handle variations in FOV and content.
@@ V-E Limitations
Small Region Segmentaiton. Figure 5 illustrates the segmentation performance of our method on challenging basal and apical slices from the ACDC dataset. Despite our method’s overall effectiveness demonstrated in previous sections, the segmentation results for these particular slices are suboptimal. This is a common challenge even for supervised methods, as the basal and apical slices often present more anatomical variations and smaller regions of interest, making accurate segmentation difficult. The queried images visualized indicate that the retrieved samples, while semantically similar, may not provide adequate support due to the small size of the target regions and the inherent heterogeneity in medical images. To quantitatively investigate this influence, we computed the DSC separately for regions smaller than 200 pixels and those larger than 200 pixels in ACDC dataset for each class. We found that segmentation performance of our method varies depending on the size of the target region. For the right ventricle (RV), the DSC is 0.2723 for small regions (<200 pixels) and increases to 0.8289 for large regions ($\geqslant$200 pixels). Similarly, for the myocardium (Myo), the DSC is 0.2226 for small regions and improves significantly to 0.8148 for large regions. For the left ventricle (LV), the DSC is 0.3402 for small regions and rises to 0.8814 for large regions. These findings align with the qualitative results shown Figure 5. This suggests that the retrieval process in our method, which relies on whole-image features, may not always capture the fine details required for accurate segmentation of small objects, which needs future efforts.